a circle with area 60 cm^2 has an arc 8 cm long the angle that is subtented at the centre of the circle by the arc is ?

Grade:12th pass

1 Answers

Arun
6 years ago
If area is given then
\pi* r² = 60 where \pi = 3.14157
Now you can find r
Now
We know the formula
L = r *\theta
Where L = length of Arc
r = radius of circle
\theta = angle made by arc at the centre
Now use this and get the answer